OpenStudy (anonymous):

Find\[ y = f(x) \]Such that \[ \int_0^tf'(x)~dy = 2t^2 \]

OpenStudy (anonymous):

I want to see what you guys come up with for this one.

OpenStudy (anonymous):

I know one solution, I'm not sure how many there are.

OpenStudy (anonymous):

Why did you delete your reply?

OpenStudy (freckles):

found my mistake

OpenStudy (anonymous):

Oh

OpenStudy (freckles):

\[\int\limits_{0}^{t}(f'(x))^2 dx=G(t)-G(0) \\ \text{ where } G(t)-G(0)=2t^2 \\ G'(t)=4t \\ \text{ but } G'(x)=(f'(x))^2 \\ g(x)=(f'(x))^2 \\ 4x=(f'(x))^2 \\ 2 \sqrt{x}=f'(x)\]

OpenStudy (freckles):

still something wrong

OpenStudy (anonymous):

Yep, something weird is going on here, lol

OpenStudy (freckles):

but this isn't too far off actually

OpenStudy (freckles):

it is a constant multiple off

OpenStudy (freckles):

oops check the wrong function

OpenStudy (freckles):

lol

OpenStudy (freckles):

I wonder what's bad about what I wrote.

OpenStudy (freckles):

oh wait it does work

OpenStudy (freckles):

sorry lol

OpenStudy (freckles):

\[f'(x)=2 \sqrt{x} \\ =2x^\frac{1}{2} \\ f(x)=2 \frac{x^\frac{3}{2}}{\frac{3}{2}}+C =2 \frac{2}{3} x^\frac{3}{2}+C=\frac{4}{3}x^\frac{3}{2}+C\]

OpenStudy (freckles):

Checking: \[f(x)=\frac{4}{3}x^\frac{3}{2}+C \\ f'(x)=2x^\frac{1}{2} \\ (f'(x))^2=4x \\ \int\limits_{0}^{t}4x dx=\frac{4x^2}{2}|_0^t=2x^2|_0^t=2(t^2-0^2)=2t^2\]

OpenStudy (freckles):

Did you get a different function that also works?

OpenStudy (anonymous):

Something not quite right is happening here. I'm not sure why though.

OpenStudy (freckles):

why my solution?

OpenStudy (freckles):

with*

OpenStudy (anonymous):

Well, consider \[ y=\frac{4}{3}x^{3/2} \]for a moment

OpenStudy (anonymous):

\[ x = \frac 34 y^{2/3} \]When we plug this into the derivative, we get: \[ 2\sqrt{x} = 2\left(\frac 34y^{2/3}\right)^{1/2} = \sqrt{3}y^{1/3} \]

OpenStudy (anonymous):

The anti derivative of that doesn't look like it is correct

OpenStudy (freckles):

\[y=\frac{4}{3}x^\frac{3}{2} \\ \frac{3}{4}y=x^\frac{3}{2} \\ (\frac{3}{4} y)^\frac{2}{3}=x \\\]

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Oh, I know what the issue is!

OpenStudy (freckles):

what?

OpenStudy (anonymous):

You made a tiny mistake

OpenStudy (anonymous):

Remember that we were integrating \(y\in [0,t]\)

OpenStudy (anonymous):

But you changed it to \(x\in [0,t]\), which is not the same region of integration.

OpenStudy (freckles):

omg

OpenStudy (freckles):

I didn't think about that

OpenStudy (freckles):

Let me try again

OpenStudy (freckles):

I forgot the limits were in terms of y :(

OpenStudy (anonymous):

Well, so far you have been very clever. Even I didn't completely understand why it wasn't working.

OpenStudy (freckles):

\[\int\limits\limits_{f^{-1}(0)}^{f^{-1}(t)}(f'(x))^2 dx=G(f^{-1}(t))-G(f^{-1}(0)) \\ \text{ where } G(f^{-1}(t))-G(f^{-1}(0))=2t^2 \\ [G(f^{-1}(t))]'=4t \\ (f^{-1})'(t)g(t)=4t \\ \frac{1}{f'(f^{-1}(t))} g(t)=4t \\ g(t)=4t f'(f^{-1}(t)) \text{ hmmm.... } \text{ we are also given } G'(x)=(f'(x))^2 \\ g(x)=(f'(x))^2 \\ g(t)=(f'(t))^2 \\ (f'(t))^2=4t f'(f^{-1}(t))\] so far I have this...

OpenStudy (freckles):

I assume f was one to one for the function I'm looking for

OpenStudy (freckles):

not successful yet still thinking

OpenStudy (freckles):

oh no I messed up lol

OpenStudy (anonymous):

To make it easier, maybe let \(x=g(y)\).

OpenStudy (anonymous):

To keep your inverses less hard to type

OpenStudy (freckles):

\[\int\limits\limits\limits_{f^{-1}(0)}^{f^{-1}(t)}(f'(x))^2 dx=G(f^{-1}(t))-G(f^{-1}(0)) \\ \text{ where } G(f^{-1}(t))-G(f^{-1}(0))=2t^2 \\ [G(f^{-1}(t))]'=4t \\ (f^{-1})'(t)g(f^{-1}(t))=4t \\ \frac{1}{f'(f^{-1}(t))} g((f^{-1}(t))=4t \\ g(f^{-1}(t))=4t f'(f^{-1}(t)) \text{ hmmm.... } \text{ we are also given } G'(x)=(f'(x))^2 \\ g(x)=(f'(x))^2 \\ \] I think I corrected it \[g(x)=4f(x)f'(x)\] \[4f(x)f'(x)=(f'(x))^2\]

OpenStudy (freckles):

\[4 f(x)=f'(x) \text{ assume } f'(x) \neq 0\]

OpenStudy (freckles):

\[4 f=\frac{df}{dx} \\ 4 dx=\frac{df}{f} \\ 4x=\ln|f(x)|+C\]

OpenStudy (freckles):

\[4x+K=\ln|f(x)| \\ e^{4x+K}=f(x) \\ f(x)=ce^{4x}\]

OpenStudy (freckles):

I have to check that

OpenStudy (anonymous):

Why, I don't understand some of your work above.

OpenStudy (anonymous):

I mean, what happened to the square around \(f'\)?

OpenStudy (freckles):

divided both sides by f'

OpenStudy (anonymous):

How did you get \(g=4ff'\)?

OpenStudy (freckles):

ok let me go through it

OpenStudy (freckles):

do you the line I said hmmm... in?

OpenStudy (anonymous):

You are correct

OpenStudy (anonymous):

I just was confused about the \(g=4ff'\) part.

OpenStudy (freckles):

\[\int\limits\limits\limits\limits_{f^{-1}(0)}^{f^{-1}(t)}(f'(x))^2 dx=G(f^{-1}(t))-G(f^{-1}(0)) \\ \text{ where } G(f^{-1}(t))-G(f^{-1}(0))=2t^2 \\ [G(f^{-1}(t))]'=4t \\ (f^{-1})'(t)g(f^{-1}(t))=4t \\ \frac{1}{f'(f^{-1}(t))} g((f^{-1}(t))=4t \\ g(f^{-1}(t))=4t f'(f^{-1}(t)) \text{ hmmm.... } \text{ we are also given } G'(x)=(f'(x))^2 \\ g(x)=(f'(x))^2 \\ \] I replaced f inverse of t with x there x=f inverse of t means f(x)=t

OpenStudy (anonymous):

So you let \(t=f(x)\)?

OpenStudy (freckles):

oh yes lol

OpenStudy (freckles):

I could have just said that

OpenStudy (anonymous):

Oh, good thinking, I did something a bit different.

OpenStudy (freckles):

did you get the same function though?

OpenStudy (anonymous):

\[ \frac{d}{dt}\int_{f^{-1}(0)}^{f^{-1}(t)}[f'(x)]^2dx = \bigg[ f'(f^{-1}(t))\bigg]^2\frac{d}{dt}f^{-1}(t) \]Then I factored out a term: \[ \bigg[ f'(f^{-1}(t))\bigg]\bigg[ f'(f^{-1}(t))\frac{d}{dt}f^{-1}(t)\bigg] \]Then here I undo the chain rule: \[ \bigg[ f'(f^{-1}(t))\bigg]\bigg[ f(f^{-1}(t))\bigg]' = \bigg[ f'(f^{-1}(t))\bigg](t)' = f'(f^{-1}(t)) \]So I had \[ f'(f^{-1}(t))=4t \]Then I plug in \(t=f(x)\)\[ f'(x) = 4f(x) \] So yes, you get correct answer.

OpenStudy (anonymous):

The correct answer is \[ y=Ce^{4x} \]I believe.

OpenStudy (freckles):

yes yes

OpenStudy (freckles):

but you were wondering if someone could get a different solution

OpenStudy (freckles):

so I wonder if we could have approached it without the assumption it was one to one

OpenStudy (freckles):

and made a different assumption on f

OpenStudy (anonymous):

What do you mean, one to one?

OpenStudy (freckles):

injective

OpenStudy (anonymous):

You mean that \(f\) is injective?

OpenStudy (freckles):

yeah

OpenStudy (freckles):

i has to be 1 to 1 (injective) to have an inverse function

OpenStudy (freckles):

f

OpenStudy (anonymous):

What if we let it equal \(f(t)\)? What is the most simplified form?

OpenStudy (freckles):

what equal f(t)?

OpenStudy (anonymous):

Find\[ y = f(x) \]Such that \[ \int_0^tf'(x)~dy = g(t) \]

OpenStudy (anonymous):

\[ f'(f^{-1}(t))=g'(t)) \]Do we stop here?

OpenStudy (anonymous):

Or maybe: \[ f'(x) = g'(f(x)) \]

OpenStudy (anonymous):

For our case, \(g(t) = 2t^2\).

OpenStudy (anonymous):

And \(g'(f(x)) = 4f(x)\).

OpenStudy (freckles):

I like this question by the way. Where did you get it?

OpenStudy (anonymous):

I made it up.

OpenStudy (freckles):

lol Also I will probably try to see if I can come up with different solution that works... But I do have to go for tonight unfortunately. P.S. So you do think there are other different solutions maybe right?

OpenStudy (anonymous):

I wonder what the identity is... if we let \(g(t)=f(t)\), then: \[ f'(x) = f'(f(x)) \]How to solve.... Maybe multiply both by \(f'(x)\): \[ [f'(x)]^2 = f'(f(x))f'(x) = [f(f(x))]' \]Woah!

OpenStudy (anonymous):

@freckles I think that your method got every possible solution. I don't think you did operations which would remove answer.

OpenStudy (freckles):

So the function f(x) doesn't have to be a function with a inverse function to write the limits as I did if that makes sense?

OpenStudy (freckles):

I mean what if f(x) has an inverse relation instead that wouldn't make a difference in the way I wrote the limits?

OpenStudy (anonymous):

Hmmm, good point. To be honest, I don't think you had to ever change the integration.

OpenStudy (anonymous):

\[ \frac{d}{dt}\int_0^tf'(x)~dy = \frac{d}{dt}\int_0^tf'(f^{-1}(y))~dy = f'(f^{-1}(t)) \]

OpenStudy (freckles):

that is the only thing I was concerned about

OpenStudy (freckles):

lol I used present tense and past tense in the same sentence I'm glad there are no language snobs around

OpenStudy (anonymous):

It's okay because it is a different clause.

OpenStudy (anonymous):

Like "I know what you did"

OpenStudy (freckles):

ah true lol

OpenStudy (freckles):

anyways thanks so much for asking a question I could answer it was a bit tricky which I like :)

OpenStudy (freckles):

goodnight